OpenStudy (anonymous):

tan theta=3/4 find sin theta

OpenStudy (anonymous):

Degress or radians?

OpenStudy (anonymous):

Or never mind, not important at all. You know that tan(angle) = opposite / adjacent? That means, if you think of a triangle with opposite = 3 and adjacent = 4, then you can use The Pythagorean Theorem to find the hypotenuse and from there it is easy to find sin(theta).

OpenStudy (jdoe0001):

recall your http://www.mathwarehouse.com/trigonometry/images/sohcohtoa/sohcahtoa-all.png \(\bf tan(\theta)=\cfrac{3}{4}\implies \cfrac{\textit{opposite side}}{\textit{adjacent side}}\implies \cfrac{b}{a}\\ \quad \\ sin(\theta)=\cfrac{\textit{opposite side}}{\textit{hypotenuse}}\implies \cfrac{b}{c}\\ \quad \\ \textit{how do we find "c"? well, use the pythagorean theorem}\\ \quad \\ c^2=a^2+b^2\implies c=\sqrt{a^2+b^2}\)

OpenStudy (anonymous):

Yes, exactly. Good reply